Crew 228 Operations Report October 4th

MDRS Operations Report 02-04-OCT-2021

Name of person filing report: Shannon Rupert and Crew Engineer David Laude

Reason for Report: Routine

Non-nominal systems: Director’s heater/Intern trailer heater

Action taken for non-nominal systems: Director’s needed a new thermostat, ordered. Intern trailer had rodents in the heater. Replaced blown fuse, will clean and assess fan, may replace, and will replace broken thermostat.

Generator: Nothing to report

ScienceDome Dual Split: Off

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 85.2

Minimum 66

Maximum 100

Note on solar: We are doing very well on solar alone for this time of year.

Diesel Reading – Empty, tank to be removed soon

Propane Reading, station tank – 76 %

Propane Reading, director tank— 71 %

Propane Reading, intern tank— 72 %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 5 gallons

Water (loft tank): ~420 gallons

Water Meter: 152,486.2 units

Water (static tank) – ~20 gallons

Static to Loft Pump used –yes

Water in GreenHab – 45% of full

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 175 gallons

Hab toilet tank emptied: no

Perseverance rover used: no

Hours:

Beginning charge: 100

Ending charge:

Currently charging: yes

Sojourner rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: yes

Hours: 140.0

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours:

Beginning charge: 100

Ending charge:

Currently charging: yes

Curiosity rover used: no

Hours:

Beginning charge: 100

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: nothing to report

HabCar used and why, where? Yes, to town for water

CrewCar used and why, where? Yes, to town laundry and supplies

Luna used and why, where? No

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal. Speed has been good with the new system.

Summary of suits and radios: Some radio rechargeable batteries are not holding enough charge during EVA. Started testing while transmitting today as a demanding test under load. Seeing a big difference among batteries. There are likely enough to get us through the mission. Will put poor batteries in labeled bag and report numbers needed as replacements to get all functional phones batteries. Will also test batteries found in a bag. This will take a few days to test all of them. After we tried on the two types of EVA suits upon arrival I forgot to mention that the 2-piece had superior air flow, enough for any occasion I can imagine. The 1-piece we are using have had enough air flow thus far.

Campus wide inspection, if action taken, what and why?

Summary of general operations: Satisfactory

Summary of Hab operations: Nominal. It would be nice during a remodel to add a large clock to each level. Looming over us they would help us stay on our schedules.

Summary of Outpost operations: Intern trailer is being prepped for painting.

Summary of GreenHab operations: Nominal, but water tank hose needs rubber ring gasket that Shannon will deliver to RAM.

Summary of ScienceDome operations: Nominal.

Summary of RAM operations: A couple tools used.

Summary of any observatory issues: Robotic observatory is down.

Summary of health and safety issues: All is good.

Questions, concerns, supplies needed and requests: Green Hab hose rubber gasket requested.

Operations Report – September 28th

MDRS Operations Report 25-28-SEPT-2021

Name of person filing report: Shannon Rupert and Crew Engineer David Laude

Reason for Report: Routine daily

Non-nominal systems: Autoclave

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Off

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 68.3

Minimum 47

Maximum 100

Note on solar: Nothing to report

Diesel Reading – Empty, tank to be removed soon

Propane Reading, station tank – 80%

Propane Reading, director tank— 74%

Propane Reading, intern tank— 72%

Propane Reading, generator— 80%

Ethanol Free Gasoline – 5 gallons

Water (loft tank): 41 gallons

Water Meter: 152,297.4 gallons

Water (static tank) – ~ 450 gallons

Static to Loft Pump used – yes

Water in GreenHab – gallons Unknown to me at this time. Will report after I find out how to measure it. Too late now.

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 130 gallons

Hab toilet tank emptied: no

Perseverance rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: yes

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: We will get hours on the first EVA.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): Yes

Reason for use: parked at Outpost

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 1

Notes on ATVs: One of the 350’s is still acting as a sign because the battery is dead. I will move tomorrow

HabCar used and why, where? Yes, for water and supplies

CrewCar used and why, where? Yes, to Grand Junction to pick up crew

Luna used and why, where? Yes, but only moved on campus

General notes and comments: Nothing to report

Summary of internet: All old accounts are nominal. New service still not working.

Summary of suits and radios: We tried on both styles. All of us picked the single piece unit. The two piece is too flawed. The lower helmet ring is too far forward as it causes the head to be forced against the back of head restricting movement and constant muscle tension against it. Also, the rear ring clamps are forced against the frame making them difficult to open or close. On some helmets the upper ring front down-pointing latch section is bent backward causing it not to engage lower clamp mate. That is easily fixed with pliers, but every time the helmet is set down the front rests upon that small piece. We would send photos if requested. Regarding the one piece, unit 7 had no fan when switched on. One battery connector had slipped off a battery terminal. Reattachment was firm, but since it came off, I will make it hold better tomorrow by crimping the female connector. One radio has suddenly become erratic and unusable. With an in-situ crew of 4, we have spares.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: RAM and SciDome were cleaned. Four signs built by the work party had pier blocks attached and they were deployed-two along Cow Dung Road stating where our land begins and ends, and two on the driveways to restrict unauthorized personnel from driving to the station. So far tourist reactions are mixed.

Summary of Hab operations: All but a few food items supplied to the crew. Lower deck cleaning completed.

Summary of Outpost operations: Nothing to report.

Summary of GreenHab operations:

Summary of ScienceDome operations: Building was cleaned today but still some things to do before it is complete

Summary of RAM operations: Organized and cleaned. Tools used to assemble 4 heavy warning signs.

Summary of any observatory issues: Robotic observatory is down.

Summary of health and safety issues: All crew nominal.

Questions, concerns, supplies needed and requests: A few food items to come. Body temp thermometer coming. Smoke detector for RAM requested.

Operations Report – September 15th

MDRS Complete Operations Report 14/15-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine daily

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Off as the temperatures aren’t so hot.

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 85.2
Minimum 63
Maximum 100

Note on solar: Happy with the uptick in SOC.

Diesel Reading – Empty, tank to be removed soon

Propane Reading, station tank – 80%

Propane Reading, director tank— 74%

Propane Reading, intern tank— 72%

Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (although I can’t find the gas can)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 225 gallons

Static to Loft Pump used – yes
Water in GreenHab – 50 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 210 gallons

Hab toilet tank emptied: no

Perseverance rover used: no
Hours: 219.4
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: yes
Hours: 147.9
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: no
Hours: 136.9
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 72.3
Beginning charge: 100
Ending charge: 95
Currently charging: no

Curiosity rover used: no
Hours: 144.4
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: I power washed Sojourner.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): Honda

Reason for use: power washed

Oil Added? No

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 would not start today, low battery

HabCar used and why, where? Yes, power washed

CrewCar used and why, where? No

Luna used and why, where? Yes, to go to town

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Nothing to report

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: We removed the old Ikea couch and moved in my little couch until the intended replacement can be purchased (currently out of stock).

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: I washed the windows. Does that count?

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Nothing to report

Operations Report – September 13th

MDRS Complete Operations Report 13-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine daily

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 75.3

Minimum 52

Maximum 100

Note on solar: We did not use the upper deck fan today. Apparently it uses a lot of power.

Diesel Reading – Empty, tank to be removed soon

Propane Reading, station tank – 80%

Propane Reading, director tank— 75%

Propane Reading, intern tank— 72%

Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 250 gallons

Static to Loft Pump used – yes

Water in GreenHab – 70 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 250 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes

Hours: 219.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: no

Hours:

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours: 136.9

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: 72.3

Beginning charge: 100

Ending charge: 95

Currently charging: no

Curiosity rover used: no

Hours: 144.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): Honda

Reason for use: put new battery in, currently running fine but needs air in rear tires

Oil Added? No

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: 350.2 would not start today, low battery

HabCar used and why, where? Yes, garbage and water

CrewCar used and why, where? No

Luna used and why, where? no

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? Low area outside SciDome that always has a puddle has been filled with sand

Summary of general operations: Nothing to report

Summary of Hab operations: Lounge chairs removed from upper deck, replaced by small sofa until new lounge couch is back in stock. Medicine cabinet from shower room was relocated to the main room to serve as our first aid kit for the Hab.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Filled the remaining carrot bin with soil.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: Any updates on possible care products for the rover’s plastic parts?

Operations Report – September 12th

MDRS Complete Operations Report 12-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:
Average 68.3
Minimum 30
Maximum 100
Note on solar: Nothing to report

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – under 100 gallons

Static to Loft Pump used – yes

Water in GreenHab – 90 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 150 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes
Hours: 219.4
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: yes
Hours:
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: yes
Hours: 136.9
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 72.3
Beginning charge: 100
Ending charge: 95
Currently charging: no

Curiosity rover used: yes
Hours: 144.4
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): 350.3
Reason for use: recreation
Oil Added? No
ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 0.5
Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, unhitched power washer and loaded to take garbage to town today

CrewCar used and why, where? No

Luna used and why, where? Yes, to Grand Junction for supplies

General notes and comments: Holes were dug for the new signs, however, we were unable to purchase the pier blocks needed for the completion of this project.

Summary of internet: All nominal.

Summary of suits and radios: Nothing to report

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Water tank and cleaned and repainted with a thick layer of elastomeric paint. The bottom had deteriorated and will need to be repaired. I have a plan to replace the lower inch or so with foam tubbing and then repaint.

Summary of Outpost operations: Water tank and cleaned and repainted with a thick layer of elastomeric paint. This tank in particular had a minimum coat of paint with many spots where the foam was exposed, so we are hoping this will protect it. The tank also needs the bottom inch or so redone.

Summary of GreenHab operations: The floor was power washed.

Summary of ScienceDome operations: Nothing to report. If the weather holds, later this week we will do a final coat on the lower portion to create a thicker coat of paint there.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: The cement pad in the back of the Hab cracked a long time ago but this summer the rains lifted in half and there is about a one-inch lift on one side. Not much can be done but warning paint will be applied.

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report – September 11th

MDRS Complete Operations Report 11-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 60.66
Minimum 34
Maximum 100

Note on solar: Nothing to report

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %

Propane Reading, director tank— 75%

Propane Reading, intern tank— 72%

Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 125 gallons

Static to Loft Pump used – yes

Water in GreenHab – 100 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 200 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes
Hours:
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: no
Hours:
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: yes
Hours:
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 72.3
Beginning charge: 100
Ending charge: 95
Currently charging: no

Curiosity rover used: yes
Hours:
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: I will report hours in the next report.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): 350.3

Reason for use: recreation

Oil Added? No

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 1
Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling power washer

CrewCar used and why, where? No

Luna used and why, where? Yes, to town

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Nothing to report

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: Nothing to report

Summary of Hab operations: Power washing complete. Hab looks nice. The water tank was too fragile for power washing and will need to be hand washed and repainted.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: The floor was power washed.

Summary of ScienceDome operations: Painting the ScienceDome was much easier than washing it. Painting with the elastomeric paint was completed in 7 hours, then 6 hours later two large storms passed to either side of the station and we got some rain.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: We had a serious case of heat exhaustion today. One member of the team passed out upstairs while trying to cool down. We took them to town in the AC cooled van for ice in soda and once they were cooled down and rested they felt better. But it was pretty scary!

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report – September 10th

MDRS Complete Operations Report 10-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action is taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 64.5
Minimum 36
Maximum 100

Note on solar: We continue to consume a lot of power considering how much solar we are generating. I believe the system can handle the load for another couple of days.

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons
Water Meter: n/a units
Water (static tank) – 200 gallons
Static to Loft Pump used – yes
Water in GreenHab – 120 gallons
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 200 gallons

Hab toilet tank emptied: yes

Perseverance rover used: Yes
Hours: 218.4
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: yes
Hours: 147.2
Beginning charge: 100
Ending charge: 100
Currently charging: no

Spirit rover used: yes
Hours: 136.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 72.3
Beginning charge: 100
Ending charge: 95
Currently charging: no

Curiosity rover used: yes
Hours: 143.3
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: We are strengthening the batteries by running the rovers in the evening

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): 350.3
Reason for use: recreation
Oil Added? No
ATV Fuel Used: 5 Gals
# Hours the ATVs were Used today: 3
Notes on ATVs: Nothing to report
HabCar used and why, where? Yes, hauling power washer

CrewCar used and why, where? Yes, for transport, the spare tire issue resolved

Luna used and why, where? No. DEF added.

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Nothing to report

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: New signs were mounted to posts and are awaiting the purchase of pier blocks, which we have been unable to find.

Summary of Hab operations: ¾ of the power washing is completed. The seam between the roof and the sides was really dirty and looks much nicer now. We determined that the structure has too much damage to paint at this time. Kay and Shannon will work with Robert to determine what is to be done.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: Cleaning was completed. We begin painting Saturday morning.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: I guess I should just stop complaining about the weather now. Nothing I can do about it.

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report – September 9th

MDRS Complete Operations Report 9-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: The seal on the Hab toilet appears to be leaking.

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 54.7

Minimum 23

Maximum 100

Note on solar: Kept rovers plugged in but did not use my AC this evening. Will keep monitoring.

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 300 gallons

Static to Loft Pump used – yes
Water in GreenHab – 120 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 200 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes
Hours:
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: no
Hours: 147.2
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: no
Hours: 135.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 72.3
Beginning charge: 100
Ending charge: 95
Currently charging: no

Curiosity rover used: yes
Hours: 141.3
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): none
Reason for use: n/a
Oil Added? Yes, to 1 and 2. 3 has very dirty old.
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling power washer to town. We have used 600 gallons of water so far
CrewCar used and why, where? No, the spare tire storage issue resolved

Luna used and why, where? No

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Nothing new to report.

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: Due to the failure of the town’s septic system, we are currently unable to pump the main system. It is ¾ full and I will just monitor it for now. We should be able to pump sometime in October.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Septic tank was emptied.

Summary of GreenHab operations: Nothing to report.

Summary of ScienceDome operations: We spent the entire day cleaning the exterior of the building to prep for painting. As clay is a charged particle, it clings to the exterior and is very hard to remove. This is why over time the building has taken on a pinkish-red hue. There was a steep learning curve, but we are finally getting it done. By nightfall we were about 80 % done, having used almost 600 gallons of water. Tomorrow morning we will finish the dome, then move on to cleaning the Hab while allowing the mud to dry around the ScienceDome before we begin painting. The only damage we caused was a small hole on the southwest side, which we will patch. The ScienceDome is white again. We also found some areas where the original paint job was only a single coat, so we have decided to do two coats of paint at this time.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Still too hot.

Questions, concerns, supplies needed and requests: Nothing at this time.

Operations Report 8-Sept-2021

MDRS Complete Operations Report 08-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power

Solar— SOC % Last 24 hours:

Average 62.9

Minimum 29

Maximum 100

Note on solar: Not sure why this happened. 29 is very low. The only new power use was plugging in the rovers and that happened late in the day. I will keep an eye on it today.

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %

Propane Reading, director tank— 75%

Propane Reading, intern tank— 72%

Propane Reading, generator: 80%

Ethanol Free Gasoline – 5 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons

Water Meter: n/a units

Water (static tank) – 350 gallons

Static to Loft Pump used – yes

Water in GreenHab – 140 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 200 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes

Hours: 217.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Sojourner rover used: yes

Hours: 147.2

Beginning charge: 100

Ending charge: 87

Currently charging: no

Spirit rover used: yes

Hours: 135.7

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: yes

Hours: 72.3

Beginning charge: 100

Ending charge: 95

Currently charging: no

Curiosity rover used: yes

Hours: 141.3

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: The remaining four rovers, with their new batteries, were delivered to the entrance to Cow Dung Road, at 1 pm. We returned them to the station. Opportunity overheated (a past issue) and so was finally retrieved after dark. Spirit’s ignition needs adjusting. Otherwise all are in fine condition, except that they need a conditioner on their exterior surfaces and Percy needs tags.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): All three 350’s

Reason for use: geology and transport to end of road to wait for rovers

Oil Added? No

ATV Fuel Used: 5 Gals

# Hours the ATVs were Used today: 3

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling water

CrewCar used and why, where? Yes, but the spare tire has come loose and we need to fix it before we use the car again

Luna used and why, where? Yes, to town and to retrieve rovers

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios and earpieces have been tested

Campus wide inspection, if action taken, what and why? Work party projects in this area are completed

Summary of general operations: It was a slower day. We have been working so hard and were all exhausted.

Summary of Hab operations: Lift boom and power washer were delivered for the paint work. We plan to focus on this large job for the remainder of the work party.

Summary of Outpost operations: General organization and cleanup of area was done. The door to the generator shack was repaired and rehung! Work party projects in this area are completed

Summary of GreenHab operations: General organization and cleanup of GreenHab was done. Work party projects in this area are completed.

Summary of ScienceDome operations: Lift boom and power washer were delivered for the paint work. We plan to focus on this large job for the remainder of the work party.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Still too hot.

Questions, concerns, supplies needed and requests: Can someone investigate and find out what kind of conditioner we should use on the rover’s exterior plastic parts? They are fading and I would like them not to flake. Thank you.

Operations Report – September 7th

MDRS Complete Operations Report 07-SEPT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Beginning of field season/work party daily update

Non-nominal systems: Refrigerator in intern trailer

Action is taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Operating

Solar— Nominal, providing all power
Solar— SOC % Last 24 hours:
Average 80
Minimum 59
Maximum 100
Note on solar: Nothing to report

Diesel Reading – 0 %.

Propane Reading, station tank – 80 %
Propane Reading, director tank— 75%
Propane Reading, intern tank— 72%
Propane Reading, generator: 80%

Ethanol Free Gasoline – 0 gallons (and it looks like we are missing a couple of gas cans)

Water (loft tank): 50 gallons
Water Meter: n/a units
Water (static tank) – 350 gallons
Static to Loft Pump used – yes
Water in GreenHab – 160 gallons
Water in ScienceDome: 0 gallons
Water (Outpost tank) – 75 gallons

Hab toilet tank emptied: no

Perseverance rover used: Yes
Hours:
Beginning charge: 100
Ending charge:
Currently charging: yes

Notes on rovers: The remaining four rovers, with new batteries, will be delivered to the entrance to Cow Dung Road, at 11 am on Wednesday.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): none
Reason for use: n/a
Oil Added? No
ATV Fuel Used: Not checked Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, hauling water

CrewCar used and why, where? Yes, to Loa for food and supplies

Luna used and why, where? Yes, emptied the remaining materials and put the seat back in

General notes and comments: Nothing to report

Summary of internet: All nominal.

Summary of suits and radios: Spacesuits are offsite. Radios are charged and are being tested.

Campus-wide inspection, if action is taken, what and why? Nothing to report

Summary of general operations: The tunnels have been completely covered and this year they look really well made as this team really worked on making them secure and looking good.

We completed a bunch of small projects today so we could focus on painting for the remainder of the work party.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: A wooden wall was constructed between the bathroom and bedroom in the intern trailer (I am very jealous as it looks amazing). The bedroom will be reserved for the Assistant Director. The water system was checked and is operational. The septic system was looked at and problems were found that will need to be corrected. The drainage pipes, ditch, and tank were all fixed up and look great.
Summary of GreenHab operations: Water prep still underway

Summary of ScienceDome operations: Tomorrow the lift boom (or whatever it is called) and the power water rentals arrive. We are planning to begin washing the ScienceDome as our main project. Because of the red bentonite clay having an electrical charge, you can’t just wash it off with water. You need to actually physically wash it off with soapy water and then rinse. So basically today we will be washing a building like it was a big baby!

Summary of RAM operations: This building has a lot of peeling paint. We are not going to work on it while the work party is happening, but it will need some attention soon.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Still miserably hot.

Questions, concerns, supplies needed and requests: Nothing at this time.